# X-ray Polarimetry of Accreting White Dwarfs: A Case Study of EX Hydrae

### v1

@@ Abstract
facilities: IXPEsoftware: ixpeobssim, isis
@@ 1 Introduction
Cataclysmic variables (CVs) are close binary systems consisting of a white dwarf (WD) undergoing accretion from a main-sequence companion via Roche lobe overflow (B. Warner, 1995). The accreted matter forms an equatorial disk as it settles onto the surface of the WD. If the WD has a sufficiently strong magnetic field (>30>30 MG; D. T. Wickramasinghe & L. Ferrario, 2000), no disk will form and accretion will occur through a magnetic channeled column directly from the Roche lobe. This type of magnetic CV (mCV) is known as a Polar. When the magnetic field is moderately strong ($\sim 5-30$ MG; D. T. Wickramasinghe & L. Ferrario, 2000), an intermediate situation will occur. An accretion disk will form, but the magnetic field disrupts it at inner radii and channels the material in broad “accretion curtains” to the magnetic poles of the WD. This type of mCV is known as an intermediate polar (IP).
@@
G. Matt (2004) predicted that the X-rays from Polars will be polarized from both scattering within the accretion column and reflection off the WD surface. For IPs, their weaker magnetic fields and thinner accretion columns will not have significant optical depth for scattering, so their should not be appreciable polarization in source emission. The only source of polarization should therefore be primarily from scattering off the surface, with the polarization degree acting as a direct measurement of the total reflection amplitude between both Thomson and Compton processes.
Here we report on an observation of EX Hya with the Imaging X-ray Polarimetry Explorer (IXPE) to make such a measurement. EX Hya presents an ideal case for this based on its orbital inclination $i_{\mathrm{inc}}=76.6\arcdeg$ and relative proximity $d=56.77$ pc, see Table 1, which together help to maximize the brightness of the source and polarization detectability in a short-duration observation. The rest of EX Hya’s important system parameters are given in Table 1, including optical polarization measurements.
@@
$M_{\mathrm{WD}}$ ($M_{\odot}$) $0.55\pm 0.15$ (a)
$M_{\mathrm{M}}$ ($M_{\odot}$) $0.085\pm 0.015$ (a)
@@
References. — (a) K. Beuermann et al. (2003), (b) C. A. L. Bailer-Jones et al. (2021), (c) M. Cropper (1986)
@@ 3 Polarization Measurements
The second method, utilizing an identical background treatment, was an event-based maximum likelihood analysis described in H. L. Marshall (2021a, b, 2024) and implemented as a Bayesian nested sampling (BNS) method by Ravi et al. (in prep). [^2] Software is soon to be available via GitHub. The BNS method characterizes each model using a log-evidence parameter $\ln Z$, where $Z$ is the likelihood. Model comparison between two hypothetical Models 1 and 2 using BNS involves the standard Bayesian statistical practice of calculating the difference $\Delta\ln Z_{12}=\ln Z_{1}-\ln Z_{2}$ and then computing a Bayes factor defined as
@@
Figure 1: Protractor plot of our X-ray polarization measurement (specifically PCUBE; blue contours) in the 2 – 3 keV band and optical polarization (orange) from Table 1. Contours are the 68, 90, and 99 percent confidence contours. The red dashed line is the direction of the magnetic field poles based on the optical polarization. Note that $\Pi_{\mathrm{optical}}/5$ is plotted for readability. North corresponds to 0 degrees.
@@
3.1 Spin Variable Polarization
The discussion above focused on the spin-averaged polarization signal. A varying polarization degree was predicted by G. Matt (2004) for the reflection component of the polarization in Polars due to their spin, so it is presumably possible for the reflection component of IPs to also vary.
To test for this variability, we used ixpeobssim to phase-fold the data over EX Hya’s spin period. We note, though, that EX Hya’s spin period is measurably decreasing, so we used the quadratic ephemeris from K. Beuermann & K. Reinsch (2024)
@@ 3.1 Spin Variable Polarization
There does not appear to be any concrete evidence of variability in the X-ray polarization (black histogram) due to each bin being below its associated MDP99 (red). As such, none can be considered a significant polarization detection. A longer observation of EX Hya, either with IXPE or a future soft X-ray polarization mission, should be taken to determine any polarization variability.
@@ 4 Estimating Shock Height
Assuming a simple 2D geometry that does not accout for the curvature of the WD, the radius of reflection is related to the angle of emission by $r=h\tan\theta_{h}$. This allows us to simplify Equation 5 to
@@
This shock height is fully consistent with the analysis from G. J. M. Luna et al. (2018), showing the efficacy of using X-ray polarization for this purpose. Importantly, though, our measurement is independent of any assumptions of the inner radius of the accretion. Our calculated shock height, in as much as our simplifying assumption can be be trusted, can be considered a more precise measure of the shock height. Furthermore, we can also use our shock height as an independent parameter to say, based on G. J. M. Luna et al. (2018), that the inner radius of the accretion disk is approximately $10\,R_{\mathrm{WD}}$.
@@ 5 Conclusions
Our measurements highlight the need for future work on EX Hya and polarization from accreting WDs in general. For EX Hya, the analysis shown here should be expanded into a full spectropolarimetric analysis. Combining our IXPE data with archival data used by G. J. M. Luna et al. (2018) will allow for a more precise determination of the poalrization across the entire 2–8 keV band. This will allow a more concrete determination of the shock height. Additionally, our system of equations for solving for $h_{0}$ could be improved with more detail. For example, the reflecting surface, i.e. the WD surface, is curved and over a extended area. A full 3-dimensional computation will be needed to properly compute how a ray scatters based on both its emitting height and the location on the surface it scatters at.
In terms of observations, EX Hya should be reobserved by IXPE for a longer total exposure time to better constrain the possible variability in the polarization signal. Based on Figure 3, a minimum of $\sim$150 ks exposure time in each phase bin to measure $\Pi$ in spin phases 0.25 – 0.75, corresponding to 600 ks total exposure. For the other two bins, we would need a total exposure time of $\sim$1.2 Ms, which is within nominal IXPE observation times for other objects.
At the same time, the strong soft X-ray polarization shows that accreting white dwarfs are potential high quality targets for soft X-ray polarimeters, such as the upcoming REDSoX sounding rocket mission (H. L. Marshall et al., 2017) and/or the proposed GOSoX mission (H. L. Marshall et al., 2021). Finally, other classes of accreting WDs (e.g., non-magnetic CVs, Polars, and Symbiotic Stars) should be observed with IXPE. Nevertheless, our analysis is the first to show the diagnostic potential X-ray polarimetry has to answer questions about accreting WD systems.

### v2

@@ Abstract
facilities: IXPEsoftware: ixpeobssim (F. Kislat et al., 2015; L. Baldini et al., 2022), isis (J. C. Houck & L. A. Denicola, 2000), queen-bee (S. Ravi et al., 2025a)
@@ 1 Introduction
Cataclysmic variables (CVs) are close binary systems consisting of a white dwarf (WD) undergoing accretion from a main-sequence companion via Roche lobe overflow (B. Warner, 1995). When the WD is only weakly magnetic, more generally referred to as “non-magnetic,” the accreted matter forms an equatorial disk as it settles onto the surface of the WD. If the WD has a sufficiently strong magnetic field, the accretion disk will be disrupted in two possible ways. When the WD dwarf’s spin and orbital periods are synchronized, the WD’s magnetic field will accrete directly from the Roche point onto its magnetic pole(s) (A. D. Schwope, 2025). This type of magnetic CV (mCV) is known as a Polar. If the spin and orbital periods are asynchronous, usually with $P_{\mathrm{spin}}<P_{\mathrm{orbit}}$, an intermediate situation will occur with such objects referred to as intermediate polars (IPs). In a classical IP, an accretion disk will form, but the magnetic field disrupts it at inner radii and channels the material onto the magnetic pole(s) of the WD.
@@
G. Matt (2004) predicted that the X-rays from Polars will be polarized from both scattering within the accretion column and reflection off the WD surface. For IPs, their weaker magnetic fields and thinner accretion columns will not have significant optical depth for scattering, so there should not be appreciable polarization in source emission. The only source of polarization should therefore be from scattering off the surface, with the polarization degree acting as a direct measurement of the total reflection amplitude between both Thomson and Compton processes.
Here we report on an observation of EX Hya with the Imaging X-ray Polarimetry Explorer (IXPE) to make such a measurement. EX Hya presents an ideal case for this based on its orbital inclination $i_{\mathrm{inc}}=78\arcdeg$ and relative proximity $d=56.77$ pc, see Table 1, which together help to maximize the brightness of the source and polarization detectability in a short-duration observation. The rest of EX Hya’s important system parameters are given in Table 1, including optical polarization measurements.
@@
$M_{\mathrm{WD}}$ ($M_{\odot}$) $0.788\pm 0.025$ (a)
$M_{\mathrm{M}}$ ($M_{\odot}$) $0.1074\pm 0.0047$ (a)
@@
References. — (a) K. Beuermann et al. (2003, and citations therein), (b) C. A. L. Bailer-Jones et al. (2021), (c) M. Cropper (1986)
@@ 3 Polarization Measurements
The second method, utilizing an identical background treatment, was an event-based maximum likelihood analysis described in H. L. Marshall (2021a, b, 2024) and implemented as a Bayesian nested sampling (BNS) method in the queen-bee software by S. Ravi et al. (2025a, b). [^2] Software is available via GitHub: https://github.com/swati-ravi/QUEEN-BEE The BNS method characterizes each model using a log-evidence parameter $\ln Z$, where $Z$ is the likelihood. Model comparison between two hypothetical Models 1 and 2 using BNS involves the standard Bayesian statistical practice of calculating the difference $\Delta\ln Z_{12}=\ln Z_{1}-\ln Z_{2}$ and then computing a Bayes factor defined as
@@
Figure 1: Protractor plot of our X-ray polarization measurement (specifically PCUBE; blue contours) in the 2 – 3 keV band. Contours are the 68, 90, and 99 percent confidence contours. Grey shaded region is the confidence in the optical polarization. The red shaded region is the direction of the magnetic field pole based on the optical polarization.
The optical polarization measurement by M. Cropper (1986) was taken over a 200 min exposure, so it shows the average position of the magnetic field over both the spin and orbital periods. The same is true for the X-ray polarization angle since the IXPE observation was taken over a total of 8 days. Thus the discussion that follows concerns the spin- and orbital-average behavior of the accretion geometry.
@@
It is worth noting the simplifications that we have used in making Figure 2. The accretion geometry in EX Hya is better described as being a “curtain” that stretches over many longitudes (J. Echevarría et al., 2016; D. V. Bisikalo et al., 2020). The asymmetry of this curtain should produce a variable polarization signal based on the (vertical) cross-sectional shape we view at any given spin and orbital phase. A spin/orbital periodic polarization degree was also predicted by G. Matt (2004) for the reflection component of the polarization in Polars due to their spin, so we next look for possible variability in the X-ray polarization.
To do so, we used ixpeobssim to phase-fold the data over EX Hya’s spin period. We note, though, that EX Hya’s spin period is measurably decreasing, so we used the quadratic ephemeris from K. Beuermann & K. Reinsch (2024)
@@
There does not appear to be any concrete evidence of variability in the X-ray polarization (black histogram) due to each bin being below its associated MDP99 (red). As such, none can be considered a significant polarization detection. Consequently, we can not consider the more detailed geometry of the accretion curtain. We are limited to only consider the spin- and orbit-averaged cross-sectional shape that the curtain will have based on our line of sight. This will be that of a cylindrical column as we show in Figure 2.
We also tested for linear polarization rotation, i.e., where the polarization degree is constant but its angle rotates in the QUQ-U plane. queen-bee includes a model for testing linear polarization rotation with time (S. Ravi et al., 2025b). The results for the 2–3 keV bin show weak preference for polarization rotation of $54_{-5}^{+4}$ deg day-1 with a Bayes Factor of $B=2.38_{-3.84}^{+1.48}$ compared with the constant polarization model. This is only a weak detection, though, so we can not confidently say the rotation is detected. A longer observation of EX Hya, either with IXPE or a future soft X-ray polarization mission will be needed to determine any polarization variability in degree and/or angle.
@@ 4 Estimating Shock Height
Assuming a simple 2D geometry that does not account for the curvature of the WD, the radius of reflection is related to the angle of emission by $r=h\tan\theta_{h}$. This allows us to simplify Equation 5 to
@@
This shock height is fully consistent with the analysis from G. J. M. Luna et al. (2018), showing the efficacy of using X-ray polarization for this purpose. Importantly, though, our measurement is independent of any assumptions of the inner radius of the accretion. Our calculated shock height, in as much as our simplifying assumption can be be trusted, can be considered a more precise measure. Furthermore, we can also use our shock height as an independent parameter to say, based on G. J. M. Luna et al. (2018), that the inner radius of the accretion disk is approximately $10\,R_{\mathrm{WD}}$.
@@ 5 Conclusions
Our measurements highlight the need for future work on EX Hya and polarization from accreting WDs in general. For EX Hya, the analysis shown here should be expanded into a full spectropolarimetric analysis. Combining our IXPE data with archival data used by G. J. M. Luna et al. (2018) will allow for an even more precise determination of the polarization across the entire 2–8 keV band. This will allow a more concrete determination of the shock height. Additionally, our system of equations for solving for $h_{0}$ could be improved with more detail. For example, the reflecting surface, i.e. the WD surface, is curved and over a extended area. A full 3-dimensional computation will be needed to properly compute how a ray scatters based on both its emitting height and the location on the surface it scatters at.
In terms of observations, EX Hya should be reobserved by IXPE for a longer total exposure time to better constrain the possible variability in the polarization signal. Based on Figure 3, a minimum of $\sim$150 ks exposure time is needed to measure $\Pi$ in spin phases 0.25 – 0.75, corresponding to 600 ks total exposure. For the other two bins, a total exposure time of $\sim$1.2 Ms is needed, which is within nominal IXPE observation times for other objects.
The strong soft X-ray polarization shows that accreting white dwarfs are potential high quality targets for soft X-ray polarimeters, such as the upcoming REDSoX sounding rocket mission (H. L. Marshall et al., 2017) and/or the proposed GOSoX mission (H. L. Marshall et al., 2021). Other classes of accreting WDs (e.g., non-magnetic CVs, Polars, and Symbiotic Stars) should be observed with X-ray polarimetric missions to determine their accretion geometries with high precision. Our analysis is the first to show the diagnostic potential X-ray polarimetry has to answer such questions about accreting WD systems.